Solve the system of equations using the elimination method. Enter the answer as an ordered pair.
x - 2y = 12
4x + 2y = 8

Found 2 solutions by stanbon, algebrahouse.com:
Answer by stanbon(75887)

Solve the system of equations using the elimination method. Enter the answer as an ordered pair.
x - 2y = 12
4x + 2y = 8
-------
Add and solve for "x":
5x = 20
x = 4
-----
Solve for "y":
x - 2y = 12
4 - 12 = 2y
y = -4
------------------
Ans: (4,-4)
-----------------------
Cheers,
Stan H.
-----------------------

Answer by algebrahouse.com(1659)

x - 2y = 12
4x + 2y = 8
------------
5x = 20 {added the two equations}
x = 4 {divided each side by 5}

x - 2y = 12 {top equations}
4 - 2y = 12 {substituted 4, in for x, into top equation}
-2y = 8 {subtracted 4 from each side}
y = -4 {divided each side by -2}

(4,-4) is the solution to the system

For more help from me, visit: www.algebrahouse.com